OpenStudy (anonymous):

ABC Is a 30 60 90 triangle if AB =36 find length BC

OpenStudy (anonymous):

which angle is 30,60,90...

OpenStudy (anonymous):

18

OpenStudy (anonymous):

please write as <A=30.<B=60...or so on...

OpenStudy (anonymous):

yep its lke that

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

then cos 60=BC/AB 1/2=BC/AB BC=36*1/2=18